# Co-MTP: A Cooperative Trajectory Prediction Framework with Multi-Temporal Fusion for Autonomous Driving

### v2

@@ I INTRODUCTION
Figure 1: The overall architecture of Co-MTP. In this framework, infrastructures share the history and their prediction results to ego CAV. Then, we construct a heterogeneous scene graph with the processed trajectory data and map information, categorizing them according to the types of objects and map elements. Next, we initialize the features of nodes and edges in the relative coordinate system of each object. The CTCA Fusion is used to update the features of the nodes and edges selected by the STSA module over K Transformer layers. Finally, we take the nodes’ hidden features from the last layer and input them into the Multimodal Decoder to obtain the multimodal trajectory prediction results.
@@ III-A Problem Definition
The input comprises two parts: history trajectories, AV’s future ground truth and map data. Consider a scenario where multiple viewpoints jointly observe a set of $Z$ objects $\Omega=\{a_{i}|i=1,2,...,Z\}$ from time step $t_{0}$ to $t_{p}$. Among them, the AV perceives $M$ objects $\Omega_{AV}=\{b_{i}|i=1,2,...,M\}$ and $b_{i}\in\Omega$, while the infrastructure perceives $N$ agents $\Omega_{infra}=\{c_{i}|i=1,2,...,N\}$ and $c_{i}\in\Omega$. All objects’ features are given, including their bounding boxes (length, width, height) and types (pedestrian, bicycle, vehicle). Define the object trajectory $i$ as $X_{i}=\left\{\bm{x}^{t_{0}}_{i},\bm{x}^{t_{1}}_{i},...,\bm{x}^{t_{h}}_{i}\right\}$ in a certain frequency, where $\bm{x}^{t}_{i}$ includes coordinates, heading, velocity, etc. Due to occlusion and other reasons, the observation could be missing in some frames. Thus, each object’s trajectory has a corresponding mask vector $\bm{m}_{i}=\left\{m^{t_{0}}_{i},m^{t_{1}}_{i},...,m^{t_{h}}_{i}\right\}$, which indicates the data validity. If the data exists at time step $t$, then $m^{t}_{i}=1$; otherwise, the value is 00. In addition, to simulate the planning trajectory from the real-world dataset, we follow the setting of [22], and use the quintic curve to interpolate the future ground truth to follow the planning mission goal. The map data includes roads and signals, and all roads are represented as polylines or polygons, such as road lines, crosswalks, and stop-lines.
@@ III-E Multimodal Decoder
where $c_{k}$ is probability of the $k-th$ mode, $\bm{y}^{t}_{k}$ is the position of the agent at the timestamp $t$, the shape of tensor $Y_{cls}$ and $Y_{reg}$ are $(N\times{K})$ and $(N\times{K}\times{T}\times{2})$. $N$ is the number of target agents, $2$ means the 2-D coordinate $(x,y)$.
@@ IV-B Compared Methods
  • PP-VIC stitch the AV and infrastructure trajectory through CBMOT, and then test it with the SOAT prediction models TNT [13] and HiVT [33].
@@
Time Dimension Model Variation Vehicle-side History Infrastructure History Fusion Method AV’s Planning Other’s Future minADE $\downarrow$ minFDE $\downarrow$ MR $\downarrow$
@@ IV-C Ablation Study
We investigate the effectiveness of multiview data processing strategies and the decoder. Ablation studies evaluate variations of Co-MTP separately in both history and future time dimensions, and the results are presented in Table II.
History Fusion. To address the two key questions of what history information should be fused and how to fuse it, we design 4 models based on three conditions: whether the vehicle-side data is available, whether the infrastructure data is available, and which fusion method to adopt. First, the infrastructure has an elevated and static position and has a better performance than the prediction with single vehicle perception. Moreover, in the table, it can be observed that as the history information from multiple views becomes more comprehensive, the prediction error gradually decreases. Additionally, comparing variants 3 and 4 reveals that the fusion method involving full infrastructure information performs better, which means the learning method can fuse the trajectory feature better from different noises and the redundant information of infrastructure can benefit the fusion. This confirms the effectiveness of our design for history information fusion.
Future Fusion. We aim to demonstrate whether future fusion, specifically the driving intentions of the AV and the prediction output of the infrastructure, are beneficial to the prediction of the AV. Similarly, we design variations for Co-MTP and analyze them alongside the baseline Co-HTTP. Overall, all models incorporating future fusion outperform those that only consider a single history time dimension. Specifically, the accuracy of 5 and 6 is lower than that of Co-MTP, indicating that both the intention of the AV and the future trajectories of other agents are crucial for the AV’s prediction task. When considering both AV planning and the other’s future trajectory from infrastructure’s prediction, the prediction performance can further be improved, as the Co-HTTP and Co-MTP show.
@@ IV-E Qualitative Result
Figure 3 present the prediction performance of Co-MTP on V2X-Seq in classic driving scenarios, such as following a vehicle, highway driving, accelerating through an intersection, and waiting to turn. In each sub-figure, the predicted trajectories closely adhere to road structure constraints and traffic safety regulations. Specifically, in (d), the target vehicle intends to turn. However, due to the influence of the object’s intention to go straight without decelerating, the target vehicle is forced to wait. This reinforces the point that future information is crucial for prediction.
@@ V CONCLUSION
To fully explore the fusion and exploit comprehensive temporal information in prediction with V2X, in this paper, we introduce the Co-MTP, the first general cooperative trajectory prediction framework with multi-temporal fusion across both history and future domains. In the history domain, we develop a heterogeneous graph to learn the fusion of the history feature from multiple agents with multi-layers Transformer to handle the issue of incomplete trajectory in history domain. Our prediction goal is to support the planning of autonomous driving, and the prediction results should reflect the potential future with a certain planning action. However, the only planning information in the future domain will make the vehicle overconfident in the planning. Thus, we further extended the heterogeneous graph to future domains to capture the future interaction among ego planning and the other vehicles’ actions from infrastructure prediction. Finally, we evaluate the Co-MTP framework on real-world dataset V2X-Seq, and the results show that Co-MTP achieves state-of-the-art performance and both history and future fusion can greatly benefit prediction. Moreover, we investigate the robustness of V2X and test the Co-MTP performance with delays and noise.

### v3

@@ I INTRODUCTION
Figure 1: The overall architecture of Co-MTP. In this framework, infrastructures share the history and their prediction results to ego AV. Then, we construct a heterogeneous scene graph with the processed trajectory data and map information, categorizing them according to the types of objects and map elements. Next, we initialize the features of nodes and edges in the relative coordinate system of each object. The CTCA Fusion is used to update the features of the nodes and edges selected by the STSA module over K Transformer layers. Finally, we take the nodes’ hidden features from the last layer and input them into the Multimodal Decoder to obtain the multimodal trajectory prediction results.
@@ III-A Problem Definition
The input comprises two parts: history trajectories, AV’s future ground truth and map data. Consider a scenario where multiple viewpoints jointly observe a set of $Z$ objects $\Omega=\{a_{i}|i=1,2,...,Z\}$ from time step $t_{0}$ to $t_{p}$. Among them, the AV perceives $M$ objects $\Omega_{AV}=\{b_{i}|i=1,2,...,M\}$ and $b_{i}\in\Omega$, while the infrastructure perceives $N$ agents $\Omega_{infra}=\{c_{i}|i=1,2,...,N\}$ and $c_{i}\in\Omega$. All objects’ features are given, including their bounding boxes (length, width, height) and types (pedestrian, bicycle, vehicle). Define the object trajectory $i$ as $X_{i}=\left\{\bm{x}^{t_{0}}_{i},\bm{x}^{t_{1}}_{i},...,\bm{x}^{t_{h}}_{i}\right\}$ in a certain frequency, where $\bm{x}^{t}_{i}$ includes coordinates, heading, velocity, etc. Due to occlusion and other reasons, the observation could be missing in some frames. Thus, each object’s trajectory has a corresponding mask vector $\bm{m}_{i}=\left\{m^{t_{0}}_{i},m^{t_{1}}_{i},...,m^{t_{h}}_{i}\right\}$, which indicates the data validity. If the data exists at time step $t$, then $m^{t}_{i}=1$; otherwise, the value is 0. In addition, to simulate the planning trajectory from the real-world dataset, we follow the setting of [22], and use the quintic curve to interpolate the future ground truth to follow the planning mission goal. The map data includes roads and signals, and all roads are represented as polylines or polygons, such as road lines, crosswalks, and stop-lines.
@@ III-E Multimodal Decoder
where $c_{k}$ is probability of the kthk-th mode, $\bm{y}^{t}_{k}$ is the position of the agent at the timestamp $t$, the shape of tensor $Y_{cls}$ and $Y_{reg}$ are $(N\times{K})$ and $(N\times{K}\times{T}\times{2})$. $N$ is the number of target agents, $2$ means the 2-D coordinate $(x,y)$.
@@ IV-B Compared Methods
  • $P$-VIC stitch the AV and infrastructure trajectory through CBMOT, and then test it with the SOAT prediction models TNT [13] and HiVT [33].
@@
Time Dimension
Model
Variation
Vehicle-side
History
Infrastructure
History
Fusion
Method
AV’s
Planning
Other’s
Future
minADE $\downarrow$ minFDE $\downarrow$ MR $\downarrow$
@@ IV-C Ablation Study
We investigate the effectiveness of multi-view data processing strategies and the decoder. Ablation studies evaluate variations of Co-MTP separately in both history and future time dimensions, and the results are presented in Table II.
History Fusion. To address the two key questions of what history information should be fused and how to fuse it, we design 4 models based on three conditions: whether the vehicle-side data is available, whether the infrastructure data is available, and which fusion method to adopt. First, the infrastructure has an elevated and static position and shows a better performance than the single-agent prediction. Moreover, as the history information from multiple views becomes more comprehensive, the prediction error gradually decreases. Furthermore, comparing variants 3 and 4 reveals that the fusion method involving full infrastructure information performs better, which means the learning method can fuse the trajectory feature better from noises, and the redundant information of infrastructure can benefit the fusion.
Future Fusion. We aim to demonstrate whether future fusion, specifically the driving intentions of the AV and the prediction output of the infrastructure, are beneficial to the prediction of the AV. Similarly, we design variations for Co-MTP and analyze them alongside the baseline Co-HTTP. Overall, all models incorporating future fusion outperform those that only consider a single history time dimension. Specifically, the accuracy of 5 and 6 is lower than that of Co-MTP, indicating that both the intention of the AV and the other’s future trajectories are crucial for the AV’s prediction task. Considering both AV planning and the other’s future trajectory from infrastructure’s prediction, the prediction performance is further improved in Co-HTTP and Co-MTP.
@@ IV-E Qualitative Result
Figure 3 presents the prediction performance of Co-MTP on V2X-Seq in classic driving scenarios, such as following, highway driving, accelerating through an intersection, and waiting to turn. In each sub-figure, the predicted trajectories closely adhere to road structure constraints and traffic safety regulations. Specifically, in (d), the target vehicle intends to turn. However, due to the influence of the object’s intention to go straight without decelerating, the target vehicle is forced to wait. This reinforces the point that future information is crucial for prediction.
@@ V CONCLUSION
To fully exploit comprehensive temporal information in prediction with V2X, we introduce the Co-MTP, the first general cooperative trajectory prediction framework with multi-temporal fusion across both history and future domains. In the history domain, we develop a heterogeneous graph to learn the history fusion from multiple agents with a multi-layer Transformer to handle the incomplete history trajectory issue. To support the future planning of autonomous driving and avoid the overconfidence of the ego vehicle in the planning, we further extended the heterogeneous graph to future domains to capture the future interaction between ego planning and the other vehicles’ actions from infrastructure prediction. Finally, we evaluate the Co-MTP framework on real-world dataset V2X-Seq, and the results show that Co-MTP achieves state-of-the-art performance and that both history and future fusion can greatly benefit prediction. Moreover, we investigate the robustness of V2X and test the Co-MTP performance with delays and noise.